OpenStudy (anonymous):

Hello my ? is c(x)=25+15x A senior citizen group is planning a bus trip. total cost of the trip is the function of total number of senior citizens that will be going. what will be the cost if 40 senior citizens go. and I cant enter units

OpenStudy (anonymous):

So wouldn't you just do c(x) = 25 + 15x set x = 40 c(x) = 25 + 15(40) then multiply and add c(x) = 625

OpenStudy (anonymous):

yea I think your right! I simply wasn't sure! I was using a different formula!

OpenStudy (anonymous):

thank you!